4x+12 (10 marks) Evaluate the integral. Use the proper notation with limits. S -da (x2+4)(x+5)
Answer #1

f 4x+12 (2+4x+5) de Partial fraction decomposition of 42+12 (22+48x+5) AxtB С + x2+4 x+5 4X+)2 G279/x+5) (^2+B)(4+5) + C/x+4)4x+12 = Ax?+ (SA+B)x+58 + C 22740 -) 4X 712 = (ATC) *2 + (SA+8)x+(58746) Comparing corresponding coefficiente On both sides HX:12=5B+YC - ② 6x5 - - 250+ SB = 20 -4C-5B =-12 - 290 = c =nd 29 A=--= -6 1-) as ( - ) +B= 4 29 29. । B=4- 40 29 76 n 29 :. A-8 29 es lats) to f de. 8 ds lx+4) 8 4x+12 (**+)(x+5) arctan ( 2 ) - 2008 B en lx+5ltc 58 + 76 58= 4 esla²+41 + 38 x 29 arctan - 29 of a delats to